OpenStudy (anonymous):

Plz help this is due today Sam organizes homes and offices for clients. She charges a $35 fee per job and then $50 per hour for her organizing services. The equation y = 50x + 35 represents the charge for her services y for x number of hours. a. Graph the equation on a separate sheet of paper. Use the graph that you create to answer parts b, c, and d. b. What is the slope of the line? What does the slope represent? c. What is the y-intercept of the line? What does the y-intercept represent? d. Natalie paid Sam $385.00 for her services. How many hours did Sam work?

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@Daniellelovee

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

help

OpenStudy (daniellelovee):

first use this  and then put those x values into the equation to get the y values

OpenStudy (anonymous):

um... i dont know how to do that they tagut me but i cant rember

OpenStudy (anonymous):

i suck at spelling

OpenStudy (daniellelovee):

you just replace x for example on the first one the x that I gave you was -2 so on the equation place -2 were x is :)

OpenStudy (texaschic101):

y = 50x + 35 In y = mx + b form, the m is the slope and the b is the y intercept....therefore, the slope of this line is 50 and the y intercept is 35. The slope (50) is just basically saying for every hr worked, it goes up 50 The y intercept (35) represents the $35 fee To graph... slope = 50 y int = (0,35) x int can be found by subbing in 0 for y y = 50x + 35 0 = 50x + 35 -50x = 35 x = - 35/50 x = - 7/10....so the x int = (-7/10,0) plot your intercepts and draw a line to connect them...use the slope of 50 if you need to. ok...he was paid 385 y = 50x + 35 385 = 50x + 35 385 - 35 = 50x 350 = 50x 350/50 = x 7 = x he worked 7 hrs

OpenStudy (daniellelovee):

^^^ no giving straight answers but GREAT job person :D

OpenStudy (texaschic101):

I tried to explain the best I could

OpenStudy (daniellelovee):

is cool I said great job ^.^